OpenStudy (anonymous):

Let f(x) = sqrt{7x^{2}+5}. Find f'(x).

OpenStudy (john_es):

You should use the chain rule, \[f(x)=\sqrt{g(x)+k}\Rightarrow f'(x)=\frac{g'(x)}{2\sqrt{g(x)+k}}\]

OpenStudy (john_es):

Do you understand it?

OpenStudy (anonymous):

Is g'(x)= 14x?

OpenStudy (john_es):

Perfect.

OpenStudy (anonymous):

so would i leave it like \[14x \div2\sqrt{7x ^{2}+5}\]?

OpenStudy (john_es):

Yes, that is the right answer.

OpenStudy (john_es):

You can now simplify a little, \[f'(x)=\frac{2x}{\sqrt{7x^2+5}}\]

OpenStudy (anonymous):

its telling me that its not the answer

OpenStudy (john_es):

Sorry, I simplified the wrong way. It must be, \[f'(x)=\frac{7x}{\sqrt{7x^2+5}}\] As you see in the last formula you wrote (only divide by 2).

OpenStudy (anonymous):

okay i was wondering how you got that 2. okay i see how you did that thank you very much!!!

OpenStudy (john_es):

;)